# definition*

### v1

@@ 1 Introduction
The first contribution of this paper is formulating maximum entropy IRL [15, 16, 2] for a diffusion model. Our formulation, named Diffusion by Maximum Entropy IRL (DxMI, pronounced "di-by-me"), is a minimax problem that jointly optimizes a diffusion model and an energy-based model (EBM). In DxMI, the EBM provides the estimated log density as the reward signal for the diffusion model. Then, the diffusion model is trained to maximize the reward from EBM while maximizing the entropy of generated samples at the same time. The maximization of entropy in DxMI facilitates exploration and stabilizes the training dynamics, as shown in reinforcement learning (RL) [17], IRL [15], and EBM [18] literature. Furthermore, the entropy maximization lets the minimax problem have an equilibrium when both the diffusion model and the EBM become the data distribution.
@@
Our second contribution is Diffusion by Dynamic Programming (DxDP), a novel maximum entropy RL algorithm for updating a diffusion model without the above-mentioned difficulties. First, DxDP mitigates the marginal entropy estimation issue by optimizing the upper bound of the original objective. In the upper bound, the marginal entropy is replaced by conditional entropies, which are easier to compute for a diffusion model. Then, DxDP removes the need for back-propagation in time by interpreting the objective as an optimal control problem and applying dynamic programming using value functions. The connection between optimal control and diffusion models is increasingly gaining attention [20, 21, 22], but DxDP is the first instance of applying discrete-time dynamic programming directly to train a diffusion model. Compared to policy gradient methods, we empirically find that DxDP converges faster and provides stronger diffusion models. As an RL algorithm, DxDP may have broader utility, such as training a diffusion model from human or AI feedback [23, 19].
@@ 3.1 Objective: Generalized Contrastive Divergence
When the models are well-specified, i.e., $p(\mathbf{x})\in\mathcal{Q}=\Pi$, Nash equilibrium of GCD is $p(\mathbf{x})=q(\mathbf{x})=\pi(\mathbf{x})$, which is identical to the solution of Eq. (4). Meanwhile, there is no need to compute the normalization constant, as the two KL divergences cancel the normalization constant out.
@@ 3.2 Alternating Update of EBM and Diffusion Model
Difficulty of Diffusion Model Update.  Ideally, diffusion model parameter $\phi$ should be updated by minimizing $KL(\pi_{\phi}||q_{\theta})=\mathbb{E}_{\pi_{\phi}(\mathbf{x})}[E(\mathbf{x})/\tau]-\mathcal{H}(\pi_{\phi}(\mathbf{x}))$. However, this update is difficult in practice for two reasons.
First, marginal entropy $\mathcal{H}(\pi_{\phi}(\mathbf{x}))$ is difficult to estimate. Discrete-time diffusion models (Eq. (1)) do not provide an efficient way to evaluate $\log\pi_{\phi}(\mathbf{x})$ required in the computation of $\mathcal{H}(\pi_{\phi}(\mathbf{x}))$, unlike some continuous-time models, e.g., continuous normalizing flows [36, 24]. Other entropy estimators based on k-nearest neighbors [37] or variational methods [38, 39] do not scale well to high-dimensional spaces.
Second, propagating the gradient through time in a diffusion model may require significant memory. Also, the gradient may explode or vanish during propagation, making the training unstable [19].
@@ 4 Diffusion by Dynamic Programming
2:  $s_{t}\leftarrow\sigma_{t}$ // AVR initialization
@@
4:     Sample $\mathbf{x}_{0:T}\sim\pi_{\phi}(\mathbf{x})$.
@@
11:        $\min_{\phi}V_{\psi}^{t+1}(\mathbf{x}_{t+1}(\phi))-\tau D\log\sigma_{t}+\frac{\tau}{2s_{t}^{2}}||\mathbf{x}_{t}-\mathbf{x}_{t+1}(\phi)||^{2}$ // $\mathbf{x}_{t+1}$ is a function of $\phi$.
@@
13:     $s_{t}\leftarrow\alpha s_{t}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$// AVR update
@@ 4.2 Dynamic Programming
In practice, a single gradient step is taken for each iteration of policy evaluation and policy improvement.
@@
Adaptive Velocity Regularization (AVR).   We additionally propose a method for systematically determining the hyperparameter $s_{t}$’s of the auxiliary distribution $\tilde{q}(\mathbf{x}_{0:T-1}|\mathbf{x}_{T})$. We can optimize $s_{t}$ such that the inequality Eq. (6) is as tight as possible by solving $\min_{s_{0},\ldots,s_{T-1}}KL(\pi_{\phi}(\mathbf{x}_{0:T})||q_{\theta}(\mathbf{x}_{T})\tilde{q}(\mathbf{x}_{0:T-1}|\mathbf{x}_{T}))$. After calculation (details in Appendix A), the optimal $s_{t}^{*}$ can be obtained analytically: $(s_{t}^{*})^{2}=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{t+1}\sim\pi}[||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}]/D$. In practice, we can use exponential moving average to compute the expectation $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{t+1}\sim\pi}[||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}]$ during training: $s_{t}\leftarrow\alpha s_{t}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$ where we set $\alpha=0.99$ for all experiment.
@@ 5 Experiments
SFT-PG [11] 10 4.32 0.584
DxMI 10 2.71 0.600
$\tau=0$ 10 3.11 0.593
Linear time cost 10 3.15 0.574
No time cost 10 5.22 0.582
DxMI + Value Guidance 10 2.63 0.603
$\tau=0$ 10 3.07 0.598
@@
5.2 Image Generation: Accelerating Diffusion Models
On image generation tasks, we show that DxMI can be used to fine-tune a diffusion model with reduced generation steps, such as $T=4$ or $10$. We test DxMI on unconditional CIFAR-10 [51] ($32\times 32$) and conditional ImageNet [52] downsampled to $64\times 64$, using three diffusion model backbones, DDPM [3], DDGAN [46], and variance exploding version of EDM [50]. The results can be found in Table 2 and 3. Starting from a publicly available checkpoint of each pretrained backbone, we first adjust the noise schedule for the target sampling steps $T$. When adjusting the noise, we follow the schedule of FastDPM [8] for DDPM and Eq. (5) of [50] for EDM. No adjustment is made for DDGAN, which is originally built for $T=4$. The adjusted models are used as the starting point of DxMI training. A single fine-tuning run takes less than 24 hours on four A100 GPUs. We set $\tau_{1}=0.1$ and $\tau_{2}=0.01$. The sigmoid time cost is used for all image generation experiments. The sample quality is measured by FID [53], Precision (Prec., [54]), and Recall (Rec., [54]). ResNet is used as our value function and is trained from scratch. More experimental details are in Appendix C.
Short-run diffusion models fine-tuned by DxMI display competitive sample quality. Unlike distillation methods, which are often limited by their teacher model’s performance, DxMI can surpass the pre-trained starting point (CIFAR-10 $T=10$). Although it does not support single-step generation, DxMI offers a principled approach to training a high-quality generative model with a moderate computation burden. Note that DDGAN does not fit the formulation of DxMI, as $\pi(\mathbf{x}_{t+1}|\mathbf{x})$ in DDGAN is not Gaussian. Nevertheless, DxMI can still enhance sample quality, showing its robustness.
@@ 5.3 Energy-Based Anomaly Detection and Localization
We use separate networks for the energy function and the value function in this experiment, as the primary goal is to obtain an accurate energy function. We employ an autoencoder architecture for the energy function, treating the reconstruction error of a sample as its energy [61]. The diffusion model and the value function are five-step time-conditioned MLPs. Unlike conventional diffusion models, DxMI allows for a flexible choice of $\pi(\mathbf{x}_{0})$. We set the initial distribution for the sampler to the data distribution applied with noise, aiming to identify the energy value more precisely near the data distribution. More experimental details can be found in C.
@@ 6 Related Work
Faster Diffusion Models.  Significant effort has been dedicated to reducing the number of generation steps in diffusion models during sampling while preserving sample quality. One popular approach is to keep the trained diffusion model unchanged and improve the sampling phase independently by tuning the noise schedule [8, 62, 63, 9], improving differential equation solvers [50, 64, 65, 66], and utilizing non-Markovian formulations [45, 67, 68]. While these methods are training-free, the sample quality can be further improved when the neural network is directly tuned for short-run sampling. Distillation methods train a faster diffusion sampler using training signal from a longer-run diffusion model, showing strong performance [10, 69, 70, 48, 71, 43, 42, 72]. However, a distilled model usually cannot outperform the teacher model. On the other hand, adversarial methods, or equivalently IRL approaches, use a discriminator to teach a diffusion model [11, 13, 46, 73, 12]. A model trained with adversarial methods may outperform a full-length diffusion model in principle, as the statistical distance between the model and the data is directly minimized. DxMI is based on the adversarial approach for training short-run diffusion models, incorporating two new ideas: maximum entropy regularization and dynamic programming.
RL for Diffusion Model.  RL is often employed to fine-tune diffusion models for a reward function. The source of the reward signal can be a computer program [23, 74, 20, 75, 76], or a human evaluator [19, 77, 25]. DxMI focuses on a setting where the estimated log data density is the reward. When RL is applied to diffusion models, the policy gradient [78] is the dominant choice [23, 74, 11, 75]. DxMI offers a value function-based approach as an alternative to the policy gradient. Maximum entropy RL for diffusion models is investigated in [20] but only in the continuous-time setting. DxDP investigates the discrete-time setting, which is more suitable for accelerating generation speed.
@@ 7 Conclusion
Limitations.  Training multiple components simultaneously, DxMI introduces several hyperparameters. To reduce the overhead of practitioners, we provide a hyperparameter exploration guideline in Appendix B. DxMI is not directly applicable to training a single-step generator. However, a diffusion model fine-tuned by DxMI can be distilled to a single-step generator.
@@
S. Yoon is supported by a KIAS Individual Grant (AP095701) via the Center for AI and Natural Sciences at Korea Institute for Advanced Study.
@@ Appendix B Guideline for Hyperparameters
13:     $s_{t}\leftarrow\alpha s_{t}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$// AVR update
@@ C.2 Image Generation
We believe the optimal $\tau$ values vary for each setting due to differences in noise schedules and magnitudes. For example, DDPM backbone is variance preserving formulation while EDM is in variance exploding formulation. Exploring a unified method for selecting the entropy regularization parameter is an interesting research direction.
@@ C.3 Anomaly Detection
Appendix D Interpretation of the policy improvement method

### v2

@@ 1 Introduction
The first contribution of this paper is formulating maximum entropy IRL [15, 16, 2] for a diffusion model. Our formulation, named Diffusion by Maximum Entropy IRL (DxMI, pronounced "di-by-me"), is a minimax problem that jointly optimizes a diffusion model and an energy-based model (EBM). In DxMI, the EBM provides the estimated log density as the reward signal for the diffusion model. Then, the diffusion model is trained to maximize the reward from EBM while simultaneously maximizing the entropy of generated samples. The maximization of entropy in DxMI facilitates exploration and stabilizes the training dynamics, as shown in reinforcement learning (RL) [17], IRL [15], and EBM [18] literature. Furthermore, the entropy maximization lets the minimax problem have an equilibrium when both the diffusion model and the EBM become the data distribution.
@@
Our second contribution is Diffusion by Dynamic Programming (DxDP), a novel maximum entropy RL algorithm for updating a diffusion model without the above-mentioned difficulties. First, DxDP mitigates the marginal entropy estimation issue by optimizing the upper bound of the original objective. In the upper bound, the marginal entropy is replaced by conditional entropies, which are easier to compute for a diffusion model. Then, DxDP removes the need for back-propagation in time by interpreting the objective as an optimal control problem and applying dynamic programming using value functions. The connection between optimal control and diffusion models is increasingly gaining attention [20, 21, 22], but DxDP is the first instance of applying discrete-time dynamic programming directly to train a diffusion model. Compared to policy gradient methods, we empirically find that DxDP converges faster and provides stronger diffusion models. As an RL algorithm, DxDP may have broader utility, such as fine-tuning a diffusion model from human or AI feedback.
@@ 3.1 Objective: Generalized Contrastive Divergence
When the models are well-specified, i.e., $p(\mathbf{x})\in\mathcal{Q}=\Pi$, Nash equilibrium of GCD is $p(\mathbf{x})=q(\mathbf{x})=\pi(\mathbf{x})$, which is identical to the solution of Eq. (4). Meanwhile, there is no need to compute the normalization constant, as the two KL divergences cancel the normalization constant out. Note that the objective function (Eq. (5)) can be negative, allowing $q(\mathbf{x})$ be closer to $p(\mathbf{x})$ than $\pi(\mathbf{x})$.
@@ 3.2 Alternating Update of EBM and Diffusion Model
Difficulty of Diffusion Model Update.  Ideally, diffusion model parameter $\phi$ should be updated by minimizing $KL(\pi_{\phi}||q_{\theta})=\mathbb{E}_{\pi_{\phi}(\mathbf{x})}[E_{\theta}(\mathbf{x})/\tau]-\mathcal{H}(\pi_{\phi}(\mathbf{x}))$. However, this update is difficult in practice for two reasons.
First, marginal entropy $\mathcal{H}(\pi_{\phi}(\mathbf{x}))$ is difficult to estimate. Discrete-time diffusion models (Eq. (1)) do not provide an efficient way to evaluate $\log\pi_{\phi}(\mathbf{x})$ required in the computation of $\mathcal{H}(\pi_{\phi}(\mathbf{x}))$, unlike some continuous-time models, e.g., continuous normalizing flows [36, 23]. Other entropy estimators based on k-nearest neighbors [37] or variational methods [38, 39] do not scale well to high-dimensional spaces. Second, propagating the gradient through time in a diffusion model may require significant memory. Also, the gradient may explode or vanish during propagation, making the training unstable [19].
@@ 4 Diffusion by Dynamic Programming
2:  $s_{t}\leftarrow\sigma_{t}$ // Initialize Adaptive Velocity Regularization (AVR)
@@
4:     Sample $\mathbf{x}_{0:T}\sim\pi_{\phi}(\mathbf{x}_{0:T})$.
@@
11:        $\min_{\phi}V_{\psi}^{t+1}(\mathbf{x}_{t+1}(\phi))+\tau\log\pi(\mathbf{x}_{t+1}|\mathbf{x}_{t})+\frac{\tau}{2s_{t}^{2}}||\mathbf{x}_{t}-\mathbf{x}_{t+1}(\phi)||^{2}$ // $\mathbf{x}_{t+1}$ is a function of $\phi$.
@@
13:     $s_{t}^{2}\leftarrow\alpha s_{t}^{2}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$// AVR update
@@ 4.2 Dynamic Programming
In practice, each iteration of policy evaluation and improvement involves a single gradient step.
@@
Adaptive Velocity Regularization (AVR).   We additionally propose a method for systematically determining the hyperparameter $s_{t}$’s of the auxiliary distribution $\tilde{q}(\mathbf{x}_{0:T-1}|\mathbf{x}_{T})$. We can optimize $s_{t}$ such that the inequality Eq. (6) is as tight as possible by solving $\min_{s_{0},\ldots,s_{T-1}}KL(\pi_{\phi}(\mathbf{x}_{0:T})||q_{\theta}(\mathbf{x}_{T})\tilde{q}(\mathbf{x}_{0:T-1}|\mathbf{x}_{T}))$. After calculation (details in Appendix A), the optimal $s_{t}^{*}$ can be obtained analytically: $(s_{t}^{*})^{2}=\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{t+1}\sim\pi}[||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}]/D$. In practice, we can use exponential moving average to compute the expectation $\mathbb{E}_{\mathbf{x}_{t},\mathbf{x}_{t+1}\sim\pi}[||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}]$ during training: $s_{t}^{2}\leftarrow\alpha s_{t}^{2}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$ where we set $\alpha=0.99$ for all experiment.
@@ 5 Experiments
SFT-PG [11] 10 4.82 0.606
DxMI 10 3.19 0.625
$\tau=0$ 10 3.77 0.613
Linear time cost 10 3.39 0.595
No time cost 10 5.18 0.595
DxMI + Value Guidance 10 3.17 0.623
$\tau=0$ 10 3.72 0.613
@@
Table 4: LSUN Bedroom $256\times 256$ unconditional image generation.
NFE FID ($\downarrow$) Prec. ($\uparrow$) Rec. ($\uparrow$)
StyleGAN2 [51] 1 2.35 0.59 0.48
Backbone: EDM
EDM [50] 79 2.44 0.71 0.67
Consistency Model [42] 2 5.22 0.68 0.39
DxMI 4 5.93 0.563 0.477
5.2 Image Generation: Training Diffusion Models with Small $T$
On image generation tasks, we show that DxMI can be used to fine-tune a diffusion model with reduced generation steps, such as $T=4$ or $10$. We test DxMI on unconditional CIFAR-10 [52] ($32\times 32$), conditional ImageNet [53] downsampled to $64\times 64$, and LSUN Bedroom [54] (256$\times$256), using three diffusion model backbones, DDPM [3], DDGAN [46], and variance exploding version of EDM [50]. The results can be found in Table 2, 3, and 4. Starting from a publicly available checkpoint of each pretrained backbone, we first adjust the noise schedule for the target sampling steps $T$. When adjusting the noise, for DDPM, we follow the schedule of FastDPM [8], and for EDM, we use Eq. (5) of [50]. No adjustment is made for DDGAN, which was originally built for $T=4$. The adjusted models are used as the starting point of DxMI training. A single CIFAR-10 run reaches the best FID in less than 4 hours on four A100 GPUs. We set $\tau_{1}=0.1$ and $\tau_{2}=0.01$. The sigmoid time cost is used for all image generation experiments. The sample quality is measured by FID [55], Precision (Prec., [56]), and Recall (Rec., [56]). ResNet is used as our value function and is trained from scratch. More experimental details are in Appendix C.2.
Short-run diffusion models fine-tuned by DxMI display competitive sample quality. Unlike distillation methods, which are often limited by their teacher model’s performance, DxMI can surpass the pre-trained starting point. Although DxMI does not support single-step generation, DxMI offers a principled approach to training a high-quality generative model with a moderate computation burden (Appendix D). Note that DDGAN does not fit the formulation of DxMI, as $\pi(\mathbf{x}_{t+1}|\mathbf{x}_{t})$ in DDGAN is not Gaussian. Nevertheless, DxMI can still enhance sample quality, showing its robustness.
@@ 5.3 Energy-Based Anomaly Detection and Localization
We use separate networks for the energy function and the value function in this experiment, as the primary goal is to obtain an accurate energy function. We employ an autoencoder architecture for the energy function, treating the reconstruction error of a sample as its energy [63]. The diffusion model and the value function are five-step time-conditioned MLPs. Unlike conventional diffusion models, DxMI allows for a flexible choice of $\pi(\mathbf{x}_{0})$. We set the initial distribution for the sampler to the data distribution applied with noise, aiming to identify the energy value more precisely near the data distribution. More experimental details can be found in Appendix C.3.
@@ 6 Related Work
Faster Diffusion Models.  Significant effort has been dedicated to reducing the number of generation steps in diffusion models during sampling while preserving sample quality. One popular approach is to keep the trained diffusion model unchanged and improve the sampling phase independently by tuning the noise schedule [8, 64, 65, 9], improving differential equation solvers [50, 66, 67, 68], and utilizing non-Markovian formulations [45, 69, 70]. While these methods are training-free, the sample quality can be further improved when the neural network is directly tuned for short-run sampling. Distillation methods train a faster diffusion sampler using training signal from a longer-run diffusion model, showing strong performance [10, 71, 72, 48, 73, 43, 42, 74]. A distilled model usually cannot outperform the teacher model, but adversarial or IRL methods may exceed full-length diffusion models. Hybrid methods [13, 12] combine distillation with adversarial loss, while other methods [46, 75] apply adversarial training to each denoising step. DxMI and SFT-PG [11] rely fully on adversarial training for final samples, allowing beneficial deviations from the diffusion path and reducing statistical distance from the data.
RL for Diffusion Model.  RL is often employed to fine-tune diffusion models for a reward function. The source of the reward signal can be a computer program [24, 76, 20, 77, 78], or a human evaluator [19, 79, 25]. DxMI focuses on a setting where the estimated log data density is the reward. When RL is applied to diffusion models, the policy gradient [80] is the dominant choice [24, 76, 11, 77]. DxMI offers a value function-based approach as an alternative to the policy gradient. Maximum entropy RL for diffusion models is investigated in [20, 81, 82] but only in the continuous-time setting. DxDP investigates the discrete-time setting, which is more suitable for accelerating generation speed.
@@
Related Theoretical Analyses.  The convergence guarantees of entropy-regularized IRL are provided in [91, 92] under the assumption of a linear reward and the infinite time horizon. Their guarantees are not directly applicable to a practical instance of DxMI, mainly due to the nonlinearity of the reward function, the continuous state and action spaces, and the finite-horizon setting. Establishing the convergence guarantee for DxMI could be an important future research direction. On the other hand, theoretical analyses have been conducted on MaxEnt RL under finite state and action spaces [93], which is relevant for the discrete version of DxDP. More focused analysis on entropy regularized RL for diffusion models is provided in [94].
@@ 7 Conclusion
Limitations.  Training multiple components simultaneously, DxMI introduces several hyperparameters. To reduce the overhead of practitioners, we provide a hyperparameter exploration guideline in Appendix B. DxMI is not directly applicable to training a single-step generator. However, a diffusion model fine-tuned by DxMI can be distilled to a single-step generator. DxMI does not offer the flexibility of using a different value of generation steps $T$ during the test time. Direct theoretical analysis of DxMI is challenging since the models are built on deep neural networks. Theoretical analysis that rationalizes the empirical results will be an important direction for future work.
@@
S. Yoon is supported by a KIAS Individual Grant (AP095701) via the Center for AI and Natural Sciences at Korea Institute for Advanced Study and IITP/MSIT (RS-2023-00220628). H. Hwang and F. Park are supported in part by IITP-MSIT grant RS-2021-II212068 (SNU AI Innovation Hub), IITP-MSIT grant 2022-220480, RS-2022-II220480 (Training and Inference Methods for Goal Oriented AI Agents), MSIT(Ministry of Science, ICT), Korea, under the Global Research Support Program in the Digital Field program(RS-2024-00436680) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation), KIAT grant P0020536 (HRD Program for Industrial Innovation), SRRC NRF grant RS-2023-00208052, SNU-AIIS, SNU-IPAI, SNU-IAMD, SNU BK21+ Program in Mechanical Engineering, SNU Institute for Engineering Research, Microsoft Research Asia, and SNU Interdisciplinary Program in Artificial Intelligence. D. Kwon is partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00252516 and No. RS-2024-00408003) and the POSCO Science Fellowship of POSCO TJ Park Foundation. Y.-K. Noh was partly supported by NRF/MSIT (No.RS-2024-00421203)) and IITP/MSIT (2020-0-01373).
@@ Appendix B Guideline for Hyperparameters
13:     $s_{t}^{2}\leftarrow\alpha s_{t}^{2}+(1-\alpha)||\mathbf{x}_{t}-\mathbf{x}_{t+1}||^{2}/D$// AVR update
@@ C.2 Image Generation
Evaluation
For computing FID, Precision, and Recall scores, we used the TensorFlow-based evaluation script provided by Consistency Models [42] repository, which is based on the codebase of [47]. All the images are saved in a PNG format when fed to the evaluation script.
@@
We believe the optimal $\tau$ values vary for each setting due to differences in noise schedules and magnitudes. For example, the DDPM backbone is a variance-preserving formulation, while EDM is a variance-exploding formulation. Exploring a unified method for selecting the entropy regularization parameter is an interesting research direction.
@@ C.3 Anomaly Detection
Appendix D Implementation and Computational Complexity of DxMI
DxMI (Algorithm 1) may seem complicated, but it largely mirrors the procedure of Max Ent IRL, with the exception of the AVR update. Notably, DxDP, the Max Ent RL subroutine within DxMI, is significantly simpler than standard actor-critic RL algorithms. Unlike these algorithms—such as Soft Actor Critic (SAC) [17], which requires training both a value function and two Q-functions—DxMI only trains a single value function and no Q-functions.
DxMI does not demand excessive computation, a major concern in image generation experiments. In these experiments, the only additional component beyond the diffusion model is the EBM, which shares the same network as the value function. Additionally, the EBM used in DxMI is typically much smaller than the diffusion model, imposing minimal computational overhead. For instance, in our CIFAR-10 experiment (T=10), the EBM consists of 5M parameters, compared to 36M in the diffusion model. This is also significantly smaller than the critic networks used in GANs, such as the 158.3M parameters in BigGAN [49]. In practice, our CIFAR-10 experiment completes in under 24 hours on two A100 GPUs, while the ImageNet 64 experiment takes approximately 48 hours on four A100 GPUs.
Appendix E Interpretation of the policy improvement method